OpenStudy (anonymous):

The average (arithmetic mean) of 5 numbers is 13. If the average of 4 of these numbers is 10, what is the fifth number?

OpenStudy (anonymous):

Let \(s\) be the sum of the and \(a\) be the fifth number. We know: \[\frac{s}{4}=10\]\[\frac{s+a}{5}=13\]So you know how to solve for a?

OpenStudy (anonymous):

Let \(s\) be the sum of the first four numbers, I meant to type.

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

Nope. First... what is \(s\)?

OpenStudy (anonymous):

Can you solve for \(s\) in the first equation first?

OpenStudy (anonymous):

Then plug your result into the second equation.

OpenStudy (anonymous):

s is 40

OpenStudy (anonymous):

Okay so plug that into the second equation:\[\frac{s+a}{5} = 13 \Rightarrow \frac{(40)+a}{5} = 13\]Now can you solve for \(a\)?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

\(5 * 13 = 65\text{ and } 65 - 40 = 25\).

OpenStudy (anonymous):

Be careful with your math.

OpenStudy (anonymous):

so 25 is the answer????

OpenStudy (anonymous):

Yeah, do you understand how I got it?

OpenStudy (anonymous):

yah. thanls